OpenStudy (anonymous):

help please (x+20) over 2 = 3x

OpenStudy (hba):

\[\frac{ x+20 }{ 2 }=3x\] 1)multiply both sides by 2 2) subtract both sides by 20 3)kaboom the answer appears.

OpenStudy (anonymous):

after i multiply 2 by both sides i got 20x=6x but i feel like that's wrong

OpenStudy (hba):

No it would be x+20=6x

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so i got x=6x so then i subtract x from both sides? and i get5x?

OpenStudy (anonymous):

5x=20 x=4

OpenStudy (anonymous):

how did you get that?(:

OpenStudy (hba):

No subtract from both sides girly x=6x-20

OpenStudy (anonymous):

@luvvikrant93

OpenStudy (anonymous):

@hba im really lost :/

OpenStudy (hba):

Now we can take the 20 on the other hand and x on the other. 6x-20=x 6x-x=20 5x=20

OpenStudy (hba):

Okay lemme start over again then @sylinan Or you can tell where i should continue.

OpenStudy (anonymous):

okay im confused because we subtracted 20 from both sides and then i got x= 6x

OpenStudy (hba):

Okay Lets see this x+20=6x Let us subtract x from both sides. Now do it.

OpenStudy (anonymous):

oh okay i get it now so when you do that is when you get 5x=20

OpenStudy (hba):

Yes now divide both sides by 5

OpenStudy (anonymous):

(x+20)/2=3x x+20=6x 20=6x-x 20=5x x=20/5 x=4

OpenStudy (anonymous):

okay and then you get four thank you!(:

OpenStudy (hba):

Welcome :)

OpenStudy (anonymous):

ur method r so long